RC meeting and Run Plan, Sunday November 28th
RC Meeting Information
RC Daily Meetings are at 3:15pm, remote connection link: https://bluejeans.com/256205934
Accelerator Status
Accelerator is delivering beam to 4 Halls. Hall A adopted a beam optics that bypasses the malfunctioning Moller Quad, Overall beam is stable except a few beam unsteady times. PD mentioned that we will probably stay this way to Dec 7th. The power supply for Moller Quad will be replaced eventually. A more final decision will probably be made at the accelerator maintenance day meeting on next Monday.
We are on plan cycling through LH2 (15uA) and LD2 (8uA) target runs, there are a few half-production runs on the runplan (with 8uA on LH2 and 4 uA on LD2) already finished.
Dummy target run (15 uA) is also finished.
Overall beam is running stable, there's no major delay since the Moller Quad is resolved.
Hall A Status and News
BBCal Trigger threshold raised from 400 mV to 450 mV Skipped HCal calibration run with SBS-OFF, BB-ON configuration, due to lack of proper HV settings for BBCal trigger.
Pending and New issues
- DAC control GUI cannot read back discriminator threshold for HCal and BBCal sum trigger. This needs to be resolved, we are now using history setting to set BBCal Threshold as we need to change the threshold for a beam position check run this afternoon. see a log by Scott on this: https://logbooks.jlab.org/entry/3951516
- Occasionally we have beam trips from accelerator, hopefully these trips will become less frequent.
- GEM Layer J0 common mode drifting higher
- Some APVs on fiber 17 and fiber 19 unstable that causing DAQ hung, since beam is stable, and we are running fine, it's not urgent to troubleshoot this
- Do we need to check BBCal calibration regarding the issue from trigger rate? (8kHz at 400mV 2GeV cut)
Subsystem Updates
- HCAL:
- BBCal:
- Timing hodoscope:
- Grinch:
- GEMS:
- Beamline:
- DAQ:
- Replay analysis:
Run Plan
SBS-11 runplan